if the velocity of a car is increased by 20%, then the minimum distance in which it can be stopped increases by

Grade:7

1 Answers

Saksham Bansal
8 years ago
As if deacceleration in both cases is kept constant then as:-
s= (v^2- u^2 )/ 2a and in both cases v=0 thus
s1= -u2/2a and s2= -(1.20u)2/2a thus change in stopping distance
(s2-s1
thus increase in stopping distance= -1.44u2/2a – (-u2)/2a = -0.44u2/2a
thus distance increases by 0.44×100=44%